OpenStudy (quickstudent):

Find the surface area of a prism with a square base that is 5 cm by 5 cm and is 10 cm tall. show all work.

OpenStudy (mathmale):

Have you sketched this prism? If not, I recommend you do. Mark the dimensions of the square base, and indicate the height of each triangle that makes up the lateral sides.

OpenStudy (quickstudent):

This is what I got: multiplying the sum of the sides by the height of the prism. (5 * 4) * 10 20 * 10 = 200 200 cm^2

OpenStudy (quickstudent):

@Google

OpenStudy (quickstudent):

@Google

OpenStudy (quickstudent):

@Will.H

OpenStudy (will.h):

Surface area = 2LW + 2(L+W)H A = 2(5 x 5) + 2(5+5)x10 = 50 + 2(10)10 =50 + 200 = 250 square cm hope that helps